Courtney’s mass is 55kg. When she is in a tuck position her radius of gyration in the saggital plane is k= 0.35m and she rotates at an angular velocity of 9.9 rad/s. How fast would she rotate if she was able to bring her knees closer to her chest so reducing k to 0.28 m


1
Expert's answer
2021-05-24T12:34:28-0400

"I_1\omega_1=I_2\omega_2\to mr_1^2\omega_1=mr_2^2\omega_2\to"


"\omega_2=r_1^2\omega_1/r_2^2=0.35^2\cdot9.9/0.28^2=15.5\ (rad/s)" . Answer